√(999x1000x1001x1002+1)  please help me to find square root

Grade:7

1 Answers

Aditya Gupta
4 years ago
hello manasuni, let y= 1000, so that 999x1000x1001x1002+1= (y – 1)y(y+1)(y+2) + 1
= y(y^2 – 1)(y+2) + 1
= (y^2 – 1)(y^2+2y) + 1
= y^4+2y^3 – y^2 – 2y + 1
= (y^2+y – 1)^2
so, √(999x1000x1001x1002+1)= y^2+y – 1
= 1000000 + 999
1000999
KINDLY APPROVE :))